# The impact of abstract and object tags on image privacy classification

### v1

@@ 5.2 Analysis of tags co-occurrence
Previously, we observed that when using a large number of tags, the model’s performance was very similar for both concrete and abstract tags. To examine whether this effect arises from the direct co-occurrence of abstract and concrete tags, we compute the Jaccard index between every pair of abstract and concrete tags within each dataset. Specifically, we use feature vectors constructed for abstract $p_{25}(i,A)$ and concrete $p_{25}(i,B)$ tag types with $k=25$ tags retained for each image. For each pair of tags $t_{a}\in A^{\prime}$ and $t_{b}\in B^{\prime}$, where $|A^{\prime}|=|B^{\prime}|=1000$, the Jaccard index is defined as
$J(t_{a},t_{b})=\frac{\sum_{[i\in D_{train}]}(\textbf{I}(i,t_{a})*\textbf{I}(i,t_{b}))}{\sum_{[i\in D_{train}]}(\textbf{I}(i,t_{a})+\textbf{I}(i,t_{b})-\textbf{I}(i,t_{a})*\textbf{I}(i,t_{b}))}$
, where $\textbf{I}(i,t)=\textbf{I}_{[p_{25}(i,t)>0]}$ indicates if $t$ is detected in image $i$ from the train subset of the dataset $D_{train}$.
@@
Table 1 reports the number of abstract-concrete tag pairs that have a Jaccard index greater than $0.5$, indicating strong co-occurrence, along with the examples. For VISPR and PrivacyAlert, only around 100 concrete tags out of 1000 have a highly co-occurring abstract tag (out of 1000), suggesting that direct co-ocurrence between concrete and abstract tags is not common. For DIPA2, this number of highly co-occurring pairs is higher, 352. However, due to the small size of this dataset, 54 of these pairs involve concepts that occur only once in the entire train set. In summary, these findings indicate that while abstract and concrete tags rarely exhibit strong direct co-occurrence, when an image is described with many tags, they begin to convey similar information about image privacy (Sec 5.1).
@@ 6 Conclusion
We analysed the importance of extracting diverse content tags for image privacy classification. While previous studies suggested that abstract tags better represent images for privacy [21], we found that the tag concreteness has a substantial impact only when a small number of tags per image description are used and the annotation instructions were open (as opposed to object-guided). We demonstrated that as the number of tags increases beyond 13, the performance gap between models trained on different tag types diminishes, and they begin to perform similarly. When only a small number of tags are available, the optimal choice of tag type becomes critical and is closely related to the subjectivity of the task: for object-guided annotation, concrete tags perform best, although abstract tags remain competitive; for subjective annotation, abstract tags provide richer information and outperform concrete ones.
Considering that tag-based or interpretable classification often explains decisions with a small number of tags, we encourage future research to incorporate not only object-oriented concepts but also abstract concepts, particularly for subjective tasks, in order to capture richer and more diverse descriptions. However, since the detection of abstract tags may be more difficult [26], with the large tag budget, such tags can be successfully replaced with concrete, object-related tags without a significant impact on the performance. Furthermore, classifier performance can benefit significantly from even modest increases in the detailedness of tag descriptions, especially for tasks where labelling is not strictly object-guided.

### v2

@@ 5.2 Analysis of tags co-occurrence
We observed that when using a large number of tags, the model’s performance was very similar for both concrete and abstract tags. To examine whether this effect arises from the direct co-occurrence of abstract and concrete tags, we compute the Jaccard index between every pair of abstract and concrete tags within each dataset. Specifically, we use feature vectors constructed for abstract $p_{25}(x_{i},A)$ and concrete $p_{25}(x_{i},B)$ tag types with $k=25$ tags retained for each image. For each pair of tags $t_{a}\in A^{\prime}$ and $t_{b}\in B^{\prime}$, where $|A^{\prime}|=|B^{\prime}|=1000$, the Jaccard index is defined as
$J(t_{a},t_{b})=\frac{\displaystyle\sum_{[x_{i}\in D_{tr}]}\hskip-4.0pt(\textbf{I}(x_{i},t_{a})\textbf{I}(x_{i},t_{b}))}{\displaystyle\sum_{[x_{i}\in D_{tr}]}\hskip-7.0pt(\textbf{I}(x_{i},t_{a})+\textbf{I}(x_{i},t_{b})-\textbf{I}(x_{i},t_{a})\textbf{I}(x_{i},t_{b}))}\text{,}$ (5)
where $\textbf{I}(x_{i},t)=\textbf{I}_{[p_{25}(x_{i},t)>0]}$ indicates if $t$ is detected in the image $x_{i}$ from the train subset of the dataset $D_{tr}$.
@@
Table 2 reports the number of abstract-concrete tag pairs that have a Jaccard index greater than $0.5$, indicating strong co-occurrence, along with the examples. For VISPR and PrivacyAlert, only around 100 concrete tags out of 1000 have a highly co-occurring abstract tag (out of 1000), suggesting that direct co-occurrence between concrete and abstract tags is uncommon. For DIPA2, the number of highly co-occurring pairs is higher, 352. However, due to the small size of this dataset, 54 of these pairs involve concepts that occur only once in the entire train set. In summary, these findings indicate that while abstract and concrete tags rarely exhibit strong direct co-occurrence, when an image is described with many tags, they begin to convey similar information about image privacy.
@@ 6 Conclusion
We analysed the importance of extracting diverse content tags for image privacy classification. While a previous study suggested that abstract tags better represent images for privacy [5], we found that as the number of tags per image description increases beyond 13, the performance gap between models trained on different tag types diminishes, and they begin to perform similarly. When only a small number of tags are available, the optimal choice of tag type becomes critical and is closely related to the subjectivity of the task: for object-guided annotation, concrete tags perform best, although abstract tags remain competitive; for subjective annotation, abstract tags provide richer information and outperform concrete ones.
Considering that tag-based or interpretable classification often explains decisions with a small number of tags, future research should incorporate not only object-oriented concepts but also abstract concepts, particularly for subjective tasks. However, since the detection of abstract tags may be more difficult [8], with a large tag budget, concrete, object-related tags can replace abstract tags without a significant impact on the performance. Furthermore, classifier performance can benefit significantly from even modest increases in the detailedness of tag descriptions, especially for tasks where labelling is not strictly object-guided.